40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13276     Entry time: Wed Aug 30 19:49:33 2017     In reply to: 13252     Reply to this: 13280
Author: gautam 
Type: Update 
Category: LSC 
Subject: REFL55 demod board debugging 

Summary:

Today I tried debugging the mysterious increase in REFL55 signal levels in the DRMI configuration. I focused on the demod board, because last week, I had tried routing these signals through different channels on the whitening board, and saw the same effect. 

Based on my tests, everything on the Demod board seems to work as expected. I need to think more about what else could be happening here - specifically do a more direct test on the whitening board.

Details:

  • The demod board is a modified D990511 (marked up schematic + high-res photo to follow).
  • Initially, I tried probing the LO signal levels at various points with the board in the eurocrate itself, with the help of an extender card.
  • But this wasn't very convenient, so I pulled the board out to the office area for more testing.
  • The 55MHz LO signal going into the board is ~0dBm (measured with Agilent network analyzer)
  • I used the active probe to check the LO levels at various points along the signal chain, which mostly consists of attenuators, ERA-5SM amplifiers, and some splitters/phase rotators.
  • Everything seemed consistent with the expected levels based on "typical" numbers for gains and insertion losses cited in the datasheets for these devices.
  • I couldn't directly measure the level at the LO input to the mixer, but measuring the input to the ERA-5SM immediately before the mixer, barring problems with this amplifier, the LO input of the mixer is being driven at >17dBm which is what it wants.
  • Next, I decided to check the gain, gain imbalance and orthogonality of the demodulation.
  • For this purpose, I restored the board to the Eurocrate, reconnected the LO input to the board, and used a second Marconi at a slightly offset frequency to drive the PD input at ~0dBm.
  • Attachment #1 - The measured outputs look pretty balanced and orthogonal. The gain is consistent with an earlier measurement I made some months ago, when things were "normal". More bullets added after Rana's questions:
    • 300 MHz bandwidth oscilloscope used to acquire the data
    • I and Q outputs were from the daughter board
    • Data was acquired via ethernet data download utility
    • 20 MHz low-pass filter turned on on the Oscilloscope while downloading the data
Quote:

I did a quick check by switching the output of the REFL55 demod board to the inputs normally used by AS55 signals on the whitening board. Setting the whitening gain to +18dB for these channels had the same effect - ADC overflow galore. So looks like the whitening board isn't to blame. I will have to check the demod board out.

 


All connections have been restored untill further debugging later in the evening.

Attachment 1: REFL55_demod_check.pdf  167 kB  Uploaded Wed Aug 30 21:00:13 2017  | Hide | Hide all
REFL55_demod_check.pdf
ELOG V3.1.3-